Direction: Solve the following problems. Show your solution and use the appropriate significant figures.

2.Given the titration of acetic acid with sodium hydroxide.

a.Calculate the molarity of an acetic acid solution if 24.37 mL of this solution are needed to neutralize 35.29 mL of 0.205M sodium hydroxide.

b.Calculate the pH of the resulting solution.
1
Expert's answer
2019-10-09T07:34:59-0400

c(CH3COOH)=35.29mL*0.205M/24.37 mL =0.30M

CH3COOH+NaOH"\to" CH3COONa+H2O

c(CH3COONa )=0.3М

pKa(CH3COOH)=4.76

pH=7+0.5*4.76+0.5*lg(0.3)=9.12